OpenStudy (anonymous):

solve the 1/2 + 2/3 + 3/4 ... 99/100 i'm forget to solve it hehehe

OpenStudy (anonymous):

\[\sum_{n=1}^{99} \frac{ n }{ n+1 }\]

OpenStudy (anonymous):

\[\sum_{n=1}^{99}[\frac{ n+1 }{ n+1 } - \frac{ 1 }{ n+1 }] =\sum_{n=1}^{99}[1 - \frac{ 1 }{ n+1 }]\]

OpenStudy (anonymous):

does this help ?

OpenStudy (anonymous):

n/n+1 is n - (1/n+1) so gets nearer and nearer to n in the limit. Might as well just feed it into Wolfram.